A woman is three times as old as her daughter. six years ago, the sum of their ages was 36. find the age of the daughter.
Answer by CubeyThePenguin(3113)

x = woman's age
y = daughter's age

x = 3y
(x-6)+(y-6) = 36

Plug in the first equation into the second.

x + y = 36 + 12 = 48
4y = 48
y = 12

The daughter is 12 years old.